OpenStudy (anonymous):

Suppose you invest $2500 at an annual interest rate of 3% compounded continuously. How much will you have in the account after 7 years? Round the solution to the nearest dollar.

OpenStudy (anonymous):

ok so 2500+3x

OpenStudy (anonymous):

annual= every year

OpenStudy (anonymous):

wait other way around 2500x+3

OpenStudy (anonymous):

wait one sec @TheSmartOne

OpenStudy (tkhunny):

Compounded CONTINUOUSLY. So far, not even close. Do you have a formula for that?

OpenStudy (anonymous):

choices a $3297 b $3084 c $3592 d $3025

OpenStudy (anonymous):

well that would of been better to know lol :) I got D because i multiply 2500(0.21) and added 525 from what i got

OpenStudy (tkhunny):

Compounded Annually \(2500(1 + 0.03)^{7} = 3,074.68\) Compounded Semi-Annually \(2500(1 + 0.03/2)^{7*2} = 3,079.39\) Compounded Quarterly \(2500(1 + 0.03/4)^{7*4} = 3,081.78\) Compounded Monthly \(2500(1 + 0.03/12)^{7*12} = 3,083.39\) Compounded Weekly \(2500(1 + 0.03/52)^{7*52} = 3,084.01\) Compounded Daily \(2500(1 + 0.03/365)^{7*365} = 3,084.17\) What is the end of this process?

OpenStudy (tkhunny):

\(2500e^{0.03\cdot 7} = 3,084.20\) -- et voilà!